Find two consecutive odd integers whose sum is 56.
Answer by checkley71(8403)

let x & (x+2) be the 2 integers.
x+(x+2)=56
2x+2=56
2x=56-2
2x=54
x=54/2
x=27 answer for the smaller interger.
27+2=29 answer for the largest integer.
proof:
27+29=56
56=56